1. a pet store sold dogs and parakeets. They counted 15 heads and 42 feet. How many dogs were there?
Answer by funmath(2933)

If you are supposed to solve this using only one variable, please resubmit it. This is how you solve it using two variables.
Let # of dogs=x
Let # of parakeets=y
Since dogs and cats have 1 head:
x+y=15
Since dogs have 4 legs and parakeets have 2:
4x+2y=42
You now have two equations and two unknowns making this solvable. If you solve the first one for y and substitute it into the second equation for y, you can solve for x.
x+y=15
-x+x+y=-x+15
y=-x+15
4x+2(-x+15)=42
4x-2x+30=42
2x+30=42
2x+30-30=42-30
2x=12
2x/2=12/2
x=6
Substitute this in the first equation for x and solve for y.
6+y=15
-6+6+y=15-6
y=9
There are 6 dogs and 9 parakeets